Suppose Google’s stock put option has a strike price of $500 and are sold at $10 per share. If you buy one contract of three months Google’s put option. Calculate your returns if the price of Google shares will either be: $550, $490, or $450 after 3 months.
Answer #1

Holding period return
1.
=(MAX(500-550,0)-10)/10=-100.00%

2.
=(MAX(500-490,0)-10)/10=0.00%

3.
=(MAX(500-450,0)-10)/10=400.00%